# Activation degree thresholds and expressiveness of polynomial neural networks

### v3

@@ 2.2 Powers of non-proportional multivariate polynomials
Let $p_{1},\dotsc,p_{k}$ be given as above and fix a general line in $\mathbb{C}^{m}$ parameterized by $\ell(t)$. As $\sum_{i}p_{i}^{r}=0$, the restriction of the polynomials $p_{i}$ to this line yields the following equality of univariate polynomials $p_{1}(\ell(t)),\dotsc,p_{k}(\ell(t))$:
$(p_{1}(\ell(t)))^{r}+\dotsb+(p_{k}(\ell(t)))^{r}=0.$
As the $p_{i}$ are pairwise non-proportional and the line parameterized by $\ell(t)$ is general, the univariate polynomials $p_{i}(\ell(t))$ are also pairwise non-proportional as they share no common zeros. By Corollary 14, we conclude that
@@
We prove the contrapositive statement when $\mathbb{K}=\mathbb{C}$. Given integers $d,k$, let $\tilde{r}(k)=6{(k-1)}^{2}-6(k-1)+1$ and $r\geq\tilde{r}(k)$ be an integer. Fix pairwise non-proportional polynomials $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ with the property that the set $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ is linearly dependent over $\mathbb{C}$. Then there exist $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ such that
@@
Let $\beta_{i}$ be a $r$th root of $\alpha_{i}$, so that
@@ 2.3 Deep networks and high activation degree
Next, we show when activation degree thresholds exist for polynomial neural networks. The proof of the following theorem follows the arguments for [25, Theorem 14], except where we leverage our Proposition 16.
@@
$\dim\mathcal{V}_{\mathbf{d},r}=\mathrm{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
@@
Fix $r>\tilde{r}(\mathbf{d})=6{m}^{2}-6m+1$. The expected dimension of a neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the minimum of two numbers in (4). We may assume
$\operatorname{edim}(\mathcal{V}_{\mathbf{d},r})=d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})\leq d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}.$
When $L=1$ and $r\geq 1$, this can be verified by direct computation. If $L>1$ then $d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}$ is no less than $12\binom{2+r^{L-1}-1}{r^{L-1}}=2r^{L-1}+$ 1, while $d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})$ is no more than $(L-1)B^{2}$ with $B:=\max\{d_{1},\dots,d_{L-1}\}$.
Thus we have
$r^{L-1}<2r^{L-1}+1<(L-1)B^{2}$
which implies that
$r<(L-1)^{1/(L-1)}B^{2/(L-1)}<2B^{2}<\tilde{r}(\mathbf{d}).$
We show by induction on the number of layers that for activation degree $r$, a general fiber of $\Psi_{\mathbf{d},r}$ is equal to the set of multi-homogeneities of any point of the fiber given by Lemma 4. The dimension of $\mathcal{V}_{\mathbf{d},r}$ is then equal to the expected dimension.
In the case that $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension as the map $\Psi_{\mathbf{d},r}$ is an isomorphism. Therefore, the activation degree threshold for the architecture $(d_{0},d_{1})$ is zero and satisfies (12).
Now assume $L>1$. Fix general weights $W_{1},W_{2},\dots,W_{L}$ and consider given weights $\tilde{W}_{1},\tilde{W}_{2}\dots,\tilde{W}_{L}$ such that
@@
by $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]$ and the output of
@@
by $[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$. Since the weights $W_{i}$ are general and $d_{i}>1$, the homogeneous polynomials $p_{\mathbf{w}_{i}}$ are pairwise non-proportional (there is a non-empty Zariski open set of the space of weight matrices where the outputs $p_{\mathbf{w}_{1}},\dotsc,p_{\mathbf{w}_{d_{L-1}}}$ are non-proportional). By examining the first coordinate of the outputs of (13) we obtain the decomposition
$\begin{split}&w_{L_{11}}p_{\mathbf{w}_{1}}^{r}+{{w}_{L_{12}}{p}_{\mathbf{w}_{2}}^{r}}+\dots+w_{L_{1d_{L-1}}}p_{\mathbf{w}_{d_{L-1}}}^{r}\\ &=\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}+{\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}_{\mathbf{w}_{d_{L-1}}}^{r}\end{split}$ (14)
where $w_{L_{ij}}$ and $\tilde{w}_{L_{ij}}$ denote the $(i,j)$ entry of the weight matrices $W_{L}$ and $\tilde{W}_{L}$ respectively. Because we have taken $r>\tilde{r}$, Proposition 16 guarantees that (14) has two proportional summands. As $p_{\mathbf{w}_{i}}$ are not pairwise proportional, no two summands on the right side may be proportional. The equality implies that neither may two summands on the left side be proportional.
By permuting terms as necessary, we may assume these proportional terms to occur as the first term on both sides. We may then scale so that $p_{\mathbf{w}_{1}}=\tilde{p}_{\mathbf{w}_{1}}$ and subtract $\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}$ to obtain
$\begin{split}(w_{L_{11}}-\tilde{w}_{L_{11}})p^{r}_{\mathbf{w}_{1}}+{{w}_{L_{12}}{p}_{\mathbf{w}_{2}}^{r}}+\dots+w_{L_{1d_{L-1}}}p^{r}_{\mathbf{w}_{d_{L-1}}}\\ =\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}^{r}_{\mathbf{w}d_{L-1}}.\end{split}$ (15)
We repeatedly apply Proposition 16 to iteratively reduce the number of terms on the right side, until the right side of (15) is zero. Then the pairwise linear independence of the $p_{\mathbf{w}_{i}}$ and Proposition 16 imply that each coefficient is also zero. Hence, up to scaling and permutation, $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]=[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$ and the entries of $W_{L}$ and $\tilde{W}_{L}$ agree in the first row. This process is repeated for each row of $W_{L}$ and $\tilde{W}_{L}$ to show that they are equal (up to scaling and permutation). Therefore the fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(W_{1},W_{2},\dots,W_{L}))$ has dimension $\sum_{i=1}^{L-1}d_{i}.$
@@
Table 1 of [28] provides several examples of neurovarieties that have the expected dimension. In this next section we show that equi-width architectures have activation degree threshold one, that is, the corresponding neurovarieties always have the expected dimension when the activation degree is two or more.
@@ 3 Equi-width setting
Let $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ be an equi-width architecture such that $d:=d_{0}=d_{1}=\dotsb=d_{L}$ and let $\sigma_{i}:\mathbb{C}^{d_{i}}\to\mathbb{C}^{d_{i}}$ be the activation map given by coordinate-wise exponentiation to the $r$-th power. A polynomial neural network
$p_{\mathbf{w}}(x)=(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x)\in\mathcal{V}_{\mathbf{d},r}$
@@
If each $W_{i}$ is invertible, then the preimage of the zero vector under each $W_{i}$ and $\sigma_{i}$ consists of only the zero vector. Thus, the polynomial neural network
$p_{\mathbf{w}}(x)=(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x)$
has only the trivial zero.
If some $W_{i}$ is singular, then let $i_{0}$ be the minimal index such that $W_{i_{0}}$ is singular and let $v\in\ker W_{i_{0}}$ be a non-zero vector. For each $j<i_{0}$, the linear map $W_{j}$ is surjective, and each $\sigma_{j}$ is surjective as well. Thus, there exists non-zero $x_{\ast}\in\mathbb{C}^{d}$ such that
@@
We now provide the main result of this section, that the activation degree threshold of an equi-width architecture is one. This is done by explicitly computing a single fiber of the parameter map $\Psi_{\mathbf{d},r}$. The dimension of this fiber provides an upper bound on the generic dimension of a fiber, and so we obtain a lower bound on the neurovariety dimension. This lower bound is in fact equal to the upper bound given by the expected dimension, thereby proving the result.
@@
If $r=1$, then $\mathcal{V}_{\mathbf{d},1}$ is the space of linear functions from $\mathbb{C}^{d}$ to $\mathbb{C}^{d}$. Therefore $\dim\mathcal{V}_{\mathbf{d},1}=d^{2}$. Since
@@
Recall the parameter map (3) for polynomial neural networks. For notational convenience, we let $S_{L,d,r}$ denote the set of parameters that are mapped to $\Psi_{\mathbf{d},r}(I_{d},\dots,I_{d})$:
$S_{L,d,r}:=\left\{(W_{1},\dots,W_{L})\in\left(\mathbb{C}^{d\times d}\right)^{L}:\Psi_{\mathbf{d},r}(I_{d},\dots,I_{d})=\Psi_{\mathbf{d},r}(W_{1},\dots,W_{L})\right\}.$ (16)
@@
$\begin{split}S_{L,d,r}=\Big{\{}(W_{1},\dots,\,PDW_{L-1},D^{-r}P^{T})\in\left(\mathbb{C}^{d\times d}\right)^{L}:(W_{1},\dots,W_{L-1})\in S_{L-1,d,r},&\\ \,P\text{ is a permutation matrix and }D\text{ is an invertible diagonal matrix}\Big{\}}.\end{split}$ (17)
@@
$W_{L}\begin{pmatrix}q_{\mathbf{w}_{1}}^{r}(x)\\ \vdots\\ q_{\mathbf{w}_{d}}^{r}(x)\end{pmatrix}=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}$ (18)
@@
$\begin{pmatrix}q_{\mathbf{w}_{1}}(x)\\ \vdots\\ q_{\mathbf{w}_{d}}(x)\end{pmatrix}:=(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x).$
Since the polynomial neural network $(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ only has the trivial zero, Proposition 21 implies that $W_{L}$ is invertible. Multiplying both sides of (18) by $W_{L}^{-1}$, each $q_{\mathbf{w}_{i}}^{r}(x)$ is a linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$. For $r>1$, this is only possible if each $q_{\mathbf{w}_{i}}$ is a monomial. Indeed, no linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$ is a pure $r$-th power except for scalar multiples of these individual monomials. Thus each $q_{\mathbf{w}_{i}}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-1}}$ and each $q_{\mathbf{w}_{i}}$ is a scalar multiple of some $x_{j}^{r^{L-2}}$. It follows that $W_{L}$ is a scaled permutation matrix—it is an invertible matrix that sends a set of monomials to a set of scalar multiples of those monomials. Thus, we have shown (17).
@@
The fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}))$ then consists of the set of multi-homogeneities and has the expected dimension $(L-1)d$, which is an upper bound on the generic dimension of a fiber. It follows from the fiber-dimension theorem [39, Theorem 1.25] that
$Ld^{2}-(L-1)d$
is a lower bound for $\dim\mathcal{V}_{\mathbf{d},r}$. As $\operatorname{edim}\mathcal{V}_{\mathbf{d},r}$ is also an upper bound and agrees with the lower bound, $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$ and $\operatorname{ActThr}(\mathbf{d})=1$. ∎

### v4

@@ 2.2 Powers of non-proportional multivariate polynomials
Let $p_{1},\dots,p_{k}$ be pairwise non-proportional polynomials on $\mathbb{C}^{d}$ such that
$p_{1}^{r}+\cdots+p_{k}^{r}=0$
holds identically. Consider the line $\ell(t)=tx+(1-t)y$ in $\mathbb{C}^{d}$ that passes through the points $x$ and $y$ and is parameterized by $t$. Restricting each $p_{i}$ to this line yields the univariate polynomials $p_{1}(\ell(t)),\dots,p_{k}(\ell(t))$ in $t$. Then
$\big{(}p_{1}(\ell(t))\big{)}^{r}+\dotsb+\big{(}p_{k}(\ell(t))\big{)}^{r}=0$
holds identically in $t$.
If $x,y\in\mathbb{C}^{d}$ are general points, one has
$p_{i}(x)/p_{j}(x)\,\neq\,p_{i}(y)/p_{j}(y),\quad{i\neq j},$
which implies
$p_{i}(\ell(1))/p_{j}(\ell(1))\,\neq\,p_{i}(\ell(0))/p_{j}(\ell(0)),\quad i\neq j.$
Therefore the univariate polynomials $p_{1}(\ell(t)),\dots,p_{k}(\ell(t))$ are pairwise non-proportional, as any quotient of a pair is non-constant. By Corollary 14, we conclude that
@@
We prove the contrapositive statement when $\mathbb{K}=\mathbb{C}$. Given integers $d,k$, let $\tilde{r}(k)=6{(k-1)}^{2}-6(k-1)+1$ and $r>\tilde{r}(k)$ be an integer. Fix pairwise non-proportional polynomials $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ with the property that the set $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ is linearly dependent over $\mathbb{C}$. Then there exist $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ such that
@@
Let $\beta_{i}$ be an $r$th root of $\alpha_{i}$, so that
@@ 2.3 Deep networks and high activation degree
Next, we show when activation degree thresholds exist for polynomial neural networks. The proof of the following theorem follows the arguments for [24, Theorem 14], except where we leverage our Proposition 16.
@@
For fixed $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=0,\dotsc,L-1$), the activation degree threshold $\operatorname{ActThr}(\mathbf{d})$ exists. That is, there exists an integer $\tilde{r}(\mathbf{d})$ such that whenever $r>\tilde{r}(\mathbf{d})$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ of the polynomial neural network has the expected dimension,
$\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}=\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}=d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})$ (15)
@@
Fix $r>\tilde{r}(\mathbf{d})=6{m}^{2}-6m+1$.
Since the expected dimension of the neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the minimum of the two numbers in (7), it suffices to prove $\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}=d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1})$ for $r>\tilde{r}(\mathbf{d})$. We show this by inducting on the number of layers $L$ to prove a general fiber of $\Psi_{{\mathbb{C},\mathbf{d},r}}$ in (4) is equal to the set of multi-homogeneities of any point of the fiber. That is, we show the containment of sets in (6) is an equality.
Recall $\Psi_{{\mathbb{C},\mathbf{d},r}}$ from (4). When $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension as the map $\Psi_{{\mathbb{C},\mathbf{d},r}}$ is an isomorphism and $\Psi_{{\mathbb{C},\mathbf{d},r}}^{-1}(\Psi_{{\mathbb{C},\mathbf{d},r}}(W_{1}))=\{W_{1}\}$.
Now assume $L>1$ and suppose that a general fiber for depth L1L-1 is completely described by the multi-homogeneity in (6). Fix general weights $(W_{1},\dots,W_{L})\in\mathbb{C}^{d_{0}d_{1}}\times\cdots\times\mathbb{C}^{d_{L-1}d_{L}}$ and consider given weights $(\tilde{W}_{1},\dots,\tilde{W}_{L})\in\mathbb{C}^{d_{0}d_{1}}\times\cdots\times\mathbb{C}^{d_{L-1}d_{L}}$ such that
@@
by $\left[Q_{({1})}(x),\dots,Q_{({d_{L-1}})}(x)\right]^{T}$ and the output of
@@
by $\left[\tilde{Q}_{({1})}(x),\dots,\tilde{Q}_{({d_{L-1}})}(x)\right]^{T}$. Since the weights $W_{i}$ are general and $d_{i}>1$, the homogeneous polynomials $Q_{({i})}$ are pairwise non-proportional (there is a non-empty Zariski open set of the space of weight matrices where the polynomials $Q_{({1})},\dots,Q_{({d_{L-1}})}$ are pairwise non-proportional).
By examining the first coordinate of the outputs of (17) we obtain the decomposition
$\begin{split}&{a_{({1})}Q_{({1})}^{r}+a_{({2})}Q_{({2})}^{r}+\dots+a_{({d_{L-1}})}Q_{({d_{L-1}})}^{r}}\\ &={\tilde{a}_{({1})}\tilde{Q}_{({1})}^{r}+\tilde{a}_{({2})}\tilde{Q}_{({2})}^{r}+\dots+\tilde{a}_{({d_{L-1}})}\tilde{Q}_{({d_{L-1}})}^{r}}\end{split}$ (18)
where $a_{({j})}$ and $\tilde{a}_{({j})}$ denote the $(1,j)$ entry of the weight matrices $W_{L}$ and $\tilde{W}_{L}$ respectively. Because we have taken $r>\tilde{r}(\mathbf{d})$, Proposition 16 guarantees that (18) has two proportional summands. As $Q_{({i})}$ are pairwise non-proportional, no two summands on the left-hand side may be proportional. The equality implies that neither may two summands on the right-hand side be proportional.
By permuting terms as necessary, we may assume these proportional terms occur as the first term on both sides. We may then rescale so that $Q_{({1})}=\tilde{Q}_{({1})}$ and subtract $\tilde{a}_{({1})}\tilde{Q}_{({1})}^{r}$ to obtain
$\begin{split}\left(a_{({1})}-\tilde{a}_{({1})}\right)Q_{({1})}^{r}+a_{({2})}Q_{({2})}^{r}+\dots+a_{({d_{L-1}})}Q_{({d_{L-1}})}^{r}\\ =\tilde{a}_{({2})}\tilde{Q}_{({2})}^{r}+\dots+\tilde{a}_{({d_{L-1}})}\tilde{Q}_{({d_{L-1}})}^{r}\end{split}$ (19)
We repeatedly apply Proposition 16 to iteratively reduce the number of terms on the right side, until the right side of (19) is zero. Then, the pairwise linear independence of the $Q_{({i})}$ and Proposition 16 imply that each coefficient is also zero. Hence, up to scaling and permutation, $\left[Q_{({1})},\dots,Q_{({d_{L-1}})}\right]^{T}=\left[\tilde{Q}_{({1})},\dots,\tilde{Q}_{({d_{L-1}})}\right]^{T}$ and the entries of $W_{L}$ and $\tilde{W}_{L}$ agree in the first row. This process is repeated for each row of $W_{L}$ and $\tilde{W}_{L}$ to show that they are equal (up to scaling and permutation). Therefore, by the inductive hypothesis the fiber $\Psi_{{\mathbb{C},\mathbf{d},r}}^{-1}(\Psi_{{\mathbb{C},\mathbf{d},r}}(W_{1},W_{2},\dots,W_{L}))$ is completely described by multi-homogeneity. As a consequence, the fiber has dimension $(d_{1}+d_{2}+\cdots+d_{L-2})+d_{L-1}.$ By the Fiber-Dimension Theorem [39, Theorem 1.25, Part (ii)] and Equation (5), the dimension of $\mathcal{V}_{\mathbf{d},r}$ satisfies
$d_{1}+d_{2}+\cdots+d_{L-1}=\sum_{i=0}^{L-1}d_{i}d_{i+1}-\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}},$
and the result follows. ∎
@@
Table 1 of [27] provides several examples of neurovarieties that have the expected dimension. In this next section, we show that equi-width architectures have activation degree threshold one, that is, the corresponding neurovarieties always have the expected dimension when the activation degree is two or more.
@@ 3 Equi-width setting
Let ${\sigma_{i}}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ be the map given by coordinate-wise exponentiation to the $r$-th power. Suppose $W_{1},\dots,W_{L}$ are in $\mathbb{C}^{d\times d}$. The map $p_{\mathbf{w}}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ given by
$p_{\mathbf{w}}(x)=(W_{L}{\sigma_{L-1}}W_{L-1}\dotsb W_{2}{\sigma_{1}}W_{1})(x)$
@@
If each $W_{i}$ is invertible, then the preimage of the zero vector under each $W_{i}$ and $\sigma_{i}$} consists of only the zero vector. Thus, the map $p_{\mathbf{w}}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ has only the trivial zero.
Conversely, if some $W_{i}$ is singular, then let $i_{0}$ be the minimal index such that $W_{i_{0}}$ is singular and let $v\in\ker W_{i_{0}}$ be a non-zero vector. For each $j<i_{0}$, the linear map $W_{j}$ is surjective, and each ${\sigma_{j}}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is surjective as well. Thus, there exists non-zero $x_{\ast}\in\mathbb{C}^{d}$ such that
@@
We now provide the main result of this section: the activation degree threshold of an equi-width architecture is one.
@@
If $r=1$, then $\mathcal{V}_{\mathbf{d},1}$ is the space of linear functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$. Therefore $\operatorname{dim}{\mathcal{V}_{\mathbf{d},1}}=d^{2}$. Since
@@
Recall that the neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the Zariski closure of the image of the polynomial map $\Psi_{\mathbf{d},r}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ where $N=Ld^{2}$ and $M={d\binom{r^{L-1}+d-1}{d-1}}$. Extending the domain of $\Psi_{\mathbf{d},r}$ to $\mathbb{C}^{N}$ results in the map (4) with the Zariski closure of the image denoted by $Y$. By Equation (5), to complete the proof, it suffices to show $\operatorname{dim}{Y}=N-(L-1)d$, which we will do using the Fiber-Dimension Theorem [39, Theorem 1.25] for (24).
For notational convenience, we let $S_{L,d,r}$ denote the set of parameters that are mapped to $\Psi_{{\mathbb{C},\mathbf{d},r}}(I_{d},\dots,I_{d})$:
$S_{L,d,r}:=\left\{(W_{1},\dots,W_{L})\in\left(\mathbb{C}^{d\times d}\right)^{L}:\Psi_{{\mathbb{C},\mathbf{d},r}}(I_{d},\dots,I_{d})=\Psi_{{\mathbb{C},\mathbf{d},r}}(W_{1},\dots,W_{L})\right\}.$ (20)
@@
[…]
@@
$W_{L}\begin{pmatrix}Q_{({1})}^{r}(x)\\ \vdots\\ Q_{({d})}^{r}(x)\end{pmatrix}=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}$ (22)
@@
$\begin{pmatrix}Q_{({1})}(x)\\ \vdots\\ Q_{({d})}(x)\end{pmatrix}:=(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x).$
Since the polynomial neural network $(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ only has the trivial zero, Proposition 21 implies that $W_{L}$ is invertible. Multiplying both sides of (22) by $W_{L}^{-1}$, each $Q_{({i})}^{r}$ is a linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$. For $r>1$, this is only possible if each $Q_{({i})}$ is a monomial. Indeed, no linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$ is a pure $r$-th power except for scalar multiples of these individual monomials. Thus each $Q_{({i})}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-1}}$ and each $Q_{({i})}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-2}}$. It follows that $W_{L}$ is a scaled permutation matrix—it is an invertible matrix that sends a set of monomials to a set of scalar multiples of those monomials. Thus, we have shown (21).
@@
We have shown that the fiber $\Psi_{{\mathbb{C},\mathbf{d},r}}^{-1}(\Psi_{{\mathbb{C},\mathbf{d},r}}(I_{d},\dotsc,I_{d}))$ is $S_{L,d,r}$ and therefore
$\operatorname{dim}{\Psi_{{\mathbb{C},\mathbf{d},r}}^{-1}(\Psi_{{\mathbb{C},\mathbf{d},r}}(I_{d},\dotsc,I_{d}))}\,=\,\operatorname{dim}{S_{L,d,r}}\,=\,(L-1)d.$
It follows from the Fiber-Dimension Theorem [39, Theorem 1.25, Part (i)] that
$N-\operatorname{dim}{Y}\,\leq\,(L-1)d.$ (23)
Using these facts on the expected dimension:
$\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}\,=\,N-(L-1)d\text{ \quad and \quad}\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}\geq\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}},$
along with (23), we have
$\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}\,\leq\,\operatorname{dim}{Y}\,=\,\operatorname{dim}{\mathcal{V}_{\mathbf{d},r}}\,\leq\,\operatorname{edim}{\mathcal{V}_{\mathbf{d},r}}.$ (24)
Therefore $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$ and $\operatorname{ActThr}(\mathbf{d})=1$. ∎